OpenStudy (anonymous):

-6z+8=z+10-7z linear equations

OpenStudy (anonymous):

@Leong

OpenStudy (anonymous):

okay so i do the same as last time okay?? cuz i bad at explaining stuff

OpenStudy (anonymous):

yes you do. :)

OpenStudy (anonymous):

-6z+8=z+10-7z -6z+6z+8=10-6z 8=10-6z -10+8=10-10-6z 2=-6z -1 over 6 =z there you go :)

OpenStudy (anonymous):

thanks. ;)

OpenStudy (anonymous):

what there is no solution @Leong

OpenStudy (anonymous):

humh?

OpenStudy (anonymous):

sr wrong one

OpenStudy (anonymous):

is -1 over 3

OpenStudy (anonymous):

there is no solution i allready had the correct answer just needed help with the steps. :P

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

are you sure your right on the steps? @Leong

OpenStudy (anonymous):

okay, do you want me to show you how to do it?

OpenStudy (anonymous):

no thats fine i know the bacis just need help with the steps. :)

OpenStudy (anonymous):

you could test, if you have the right answer is no solution then you should have this -6 times (-1 over 3) +8 =10 ( -1 over 3) + 10 - 7 times (-1 over 3) =12 since 10 can't equal to 12 then it should be NO SOLUTION

OpenStudy (anonymous):

thank. :)

OpenStudy (anonymous):

:) ur welcome